The tank of a car is 5/6 full at the start of the journey. After a while, it was only 5:8 full. The meter showed 10 liters have been used. What is the capacity of the tank?
Answer by Alan3354(69443)

The tank of a car is 5/6 full at the start of the journey. After a while, it was only 5:8 full. The meter showed 10 liters have been used. What is the capacity of the tank?
------------
5:8 full?
If you mean 5/8
-----------
Tk*(5/6 - 5/8) = 10
Tk*(1/6 - 1/8) = 2
Tk*(8 - 6)/48 = 2
Tk = 48 liters